OpenStudy (anonymous):

x-3 over x+7=3/5 what is x?

OpenStudy (anonymous):

I got an answer of -22 if I did it right!?

OpenStudy (anonymous):

\[\frac{x-3}{x+7}=\frac{3}{5}\] cross multiply \[5(x-3)=3(x+7)\] \[5x-15=3x+21\] \[5x-3x=21+15\] \[2x=36\] \[x=18\]